# Pushing Trade-Off Boundaries: Compact yet Effective Remote Sensing Change Detection

### v1

@@ Abstract
Remote sensing change detection is essential for monitoring urban expansion, disaster assessment, and resource management, offering timely, accurate, and large-scale insights into dynamic landscape transformations. While deep learning has revolutionized change detection, the increasing complexity and computational demands of modern models have not necessarily translated into significant accuracy gains. Instead of following this trend, this study explores a more efficient approach, focusing on lightweight models that maintain high accuracy while minimizing resource consumption, which is an essential requirement for on-satellite processing. To this end, we propose FlickCD, which means quick flick then get great results, pushing the boundaries of the performance-resource trade-off. FlickCD introduces an Enhanced Difference Module (EDM) to amplify critical feature differences between temporal phases while suppressing irrelevant variations such as lighting and weather changes, thereby reducing computational costs in the subsequent change decoder. Additionally, the FlickCD decoder incorporates Local-Global Fusion Blocks, leveraging Shifted Window Self-Attention (SWSA) and Enhanced Global Self-Attention (EGSA) to efficiently capture semantic information at multiple scales, preserving both coarse- and fine-grained changes. Extensive experiments on four benchmark datasets demonstrate that FlickCD reduces computational and storage overheads by more than an order of magnitude while achieving state-of-the-art (SOTA) performance or incurring only a minor (<1% F1) accuracy trade-off. The implementation code is publicly available at https://github.com/xulsh8/FlickCD.
Keywords Change Detection, Remote Sensing, Lightweight Deep Learning
@@ 1 Introduction
With the rapid advancement of remote sensing technologies, it has become possible to acquire an increasing number of high-resolution remote sensing images. These abundant and high-quality data sources have greatly enhanced the development of various remote sensing applications, among which change detection (CD) has particularly benefited as a key task. The target of CD is to identify the changes in the same region by analyzing multi-temporal remote sensing images acquired at different times. This technique can be applied to various fields, such as disaster damage assessment [1, 2], urban planning [3, 4], agricultural management [5] and environment monitoring [6]. The CD techniques makes it possible to gain a more comprehensive understanding of the regional differences, enabling the prediction of future changes and the formulation of corresponding measures.
@@
With the growing success of deep learning in various computer vision tasks, it has also emerged as a powerful tool for CD. However, deep learning-based CD methods face two major challenges. The first challenge is how to better balance model performance and resource requirements. To achieve better feature extraction capabilities, change detection models often increase their depth or width a lot. For example, RCTNet [7] increases depth by stacking convolutional layers to fuse multi-scale feature maps, while ChangeMamba [8] expands the receptive field using state-space models to capture global semantic information. However, these operations require a substantial number of parameters and extensive computational resources, making it challenging to deploy them on resource-constrained devices, such as drones and satellites, which are commonly used for acquiring remote sensing images. Currently, there are also researches on lightweight change detection models [9, 10]. Nevertheless, achieving lightweight design often results in performance degradation and leads to a reduced ability to extract difference information.
@@
The second challenge is how to ignore irrelevant differences while focusing on capturing significant changes. Irrelevant differences can arise due to various factors, including lighting variations, atmospheric cloud conditions, seasonal changes and so on. To extract the critical changed information between two temporal images, various change detection models employ different approaches. Most models obtain difference information by computing the Manhattan distance between corresponding pixels in two temporal images [11, 9, 7]. Some models merge the two temporal feature maps along the channel dimension and use convolution operations to extract difference information [12, 10], while others perform layer-wise or multi-level feature fusion to achieve this [8, 13]. However, each of these methods has its own limitations: some of them fail to filter out irrelevant changes, making it difficult for the model to distinguish between insignificant differences and crucial changes. Others require substantial computational resources and a large number of parameters to obtain difference information, thereby increasing the overall resource demand of the model.
@@
To address the aforementioned challenges, we propose FlickCD, a lightweight model that achieves an excellent balance between resource efficiency and performance. As shown in Figure 1, we select several lightweight models and state-of-the-art models for comparison. It can be observed that while maintaining a lightweight design, FlickCD achieves an average performance close to that of state-of-the-art models.
FlickCD adopts an encoder-decoder architecture, with RepViT [14] chosen as its encoder. This encoder has the advantages of low latency and a small number of parameters. Unlike most change detection models, we utilize only the feature maps extracted from the first three stages of the encoder, as we observed that incorporating higher-level features does not lead to significant performance improvement, while substantially increasing the number of parameters and computational cost. After extracting feature maps from two temporal images using a shared-weight encoder, we employ the Enhanced Difference Module(EDM) to capture the difference information between them. The lightweight EDM can enhance the difference features while suppressing the irrelevant noise, thereby helping the decoder more accurately extract the changed features of interest. Then, we design the decoder using a bottom-up, layer-by-layer fusion approach. Before fusion, we incorporate a Local-Global Fusion Block(LGFB) to further integrate features. LGFB primarily utilizes attention mechanisms and consist of two components: Sliding-Window Self-attention(SWSA) and Effective Global Self-attention(EGSA). SWSA applies windows to confine attention to local areas, enabling the model to focus more on the relationship between nearby pixels and improving the accuracy of local object recognition. Subsequently, EGSA is employed to consider the relationship between pixels across the global scope, which is particularly crucial when the changed regions consist of a lot of similar small areas. After multiscale extraction and fusion, the process yields a high-quality binary change map as the final result.
Through extensive experiment, we validate the effectiveness of the modules in FlickCD and demonstrate that FlickCD achieves a better trade-off and higher efficiency compared to other models. Our contributions can be summarized as follows:
@@
  • We conducted experiments to demonstrate the efficiency of FlickCD among current change detection models. It achieves low parameter and computation costs while maintaining high performance, even achieves state-of-the-art on one dataset.
@@ 2.1 Deep Learning-based Change Detection
With the remarkable success and high performance of deep learning techniques, the change detection (CD) community has increasingly adopted them to address the CD task. One of the earliest CNN-based approaches is the FC series [15], which introduced two fundamental strategies for handling dual-temporal inputs: (1) early fusion, where the two input images are first concatenated and then fed into the network; and (2) Siamese architectures, which process the two images with shared-weight encoders followed by a difference computation. These two strategies have laid the foundation for most subsequent CD models. After that, the emergence of Transformer architectures, which leverage attention mechanisms, has brought new possibilities to CD. For instance, BIT [11] adopts a CNN backbone for encoding and decoding, while employing a Transformer to capture differences between the input images. ChangeFormer [12] further advances this direction by directly applying a shared-weight Transformer to handle the CD task. In addition to the conventional global self-attention, various attention mechanisms tailored for CD have also been explored. These include spatial and channel attention [16, 17, 18], as well as cross-attention mechanism [19], which enhance the effectiveness and robustness of CD models.
@@ 2.2 Lightweight Strategies in Deep Learning
Striking a perfect balance between model efficiency and performance is always a critical research direction. To enable the development of deep learning models on resource-constrained devices, a variety of lightweight strategies have been proposed across different model architectures. In convolutional neural networks(CNNs)-based strategies, depthwise separable convolution [20] extracts features by decoupling the spatial and channel dimensions. Group convolution [21] divides channels into groups and performing convolution operations within each group independently. More recently, FasterNet [22] introduce partial convolution, which compute only a subset of channels in the feature maps. In Transformer-based strategies, they are primarily implemented in two ways. First, limit the attention computation scope to local windows or sparse regions [14, 23, 24]. Although this strategy may result in information loss, allocating limited parameters to more likely relevant tokens can lead to minimal performance degradation while reducing redundant information. Second, utilize weight-sharing mapping to obtain the attention matrix [23, 25]. Applying the same linear mapping to different regions with correlations can reduce the number of parameters and map them to the same high-dimensional space. As for the CD tasks, there has also been research on lightweight models. USSFC-Net [9] uses dilated convolutions cyclically as the basic module, and construct the Encoder and Decoder by stacking this module. TinyCD [10] employs the three-stage EfficientNet [26] as its backbone and utilizes group convolution to compute difference information. Although these networks have a low number of parameters, they suffer from high latency and lower performance.
@@ 3.1 Overview
The overview of the model structure is shown in the left side of Figure 2. At first, input the bitemporal images, $T_{1}$ and $T_{2}$, both with dimensions $W\times H\times 3$, into the weight-sharing encoder. We selected the lightest version of RepViT [14] with pre-trained weights as the encoder. Then, we only use the feature maps from the first three stages of the encoder. On one hand, our experiments demonstrate that the feature map from the fourth stage does not improve the model’s performance. On the other hand, removing the fourth stage can reduce the number of parameters and computational complexity. This process can be expressed as:
$f_{1\_1},f_{1\_2},f_{1\_3}=Encoder(T_{1}),$
$f_{2\_1},f_{2\_2},f_{2\_3}=Encoder(T_{2}).$
The encoder extracts features hierarchically, where downsampling and channel doubling are applied at each stage. For the feature maps $f_{1\_j}$ and $f_{2\_j},j=(1,2,3)$, the spatial resolution at each stage is $\frac{W}{2^{j+1}}\times\frac{H}{2^{j+1}}$. Then, we use the Enhanced Difference Module(EDM) to extract the difference feature between $f_{1}$ and $f_{2}$ at each stage. Additionally, to facilitate the multi-stage feature fusion in the decoder, the channel dimension of the feature maps from all three stages are unified. We formulate this process as:
$D_{i}=EDM_{i}(f_{1\_i},f_{2\_i}),\{i=1,2,3\}.$
After that, we propose a bottom-up lightweight decoder, which mainly consists of the Local-Global Fusion Block(LGFB) and additional Effective Global Self-attention(EGSA). Through layer-by-layer fusion and upsampling, it generates the final binary result map.
@@ 3.2 Enhanced Difference Module
How to extract the difference information between two temporal images effectively is one of the main challenges of the CD task. A common approach for extracting the difference is to directly subtract the two feature maps and take the absolute value, which corresponds to the Manhattan distance. However, since this approach computes pixel-wise differences with equal weighting across all pixels, trivial differences caused by irrelevant factors are preserved alongside meaningful change information and passed to the decoder. This makes it difficult for the decoder to accurately identify critical differences, leading to a decrease in accuracy. Moreover, additional computational resources and parameters may be required to effectively distinguish between relevant and irrelevant information. To address this issue, we propose the Enhanced Difference Module(EDM), which dynamically assigns weights to the pixel-wise differences. This enables the model to filter out irrelevant difference and enhance the identification of critical change information.
@@
The detailed architecture of the EDM is illustrated on the right side of Fig 2. For a given stage $EDM_{i}$, the inputs are the corresponding temporal feature maps $f_{1\_i}$ and $f_{2\_i}$. Since the magnitude of pixel-wise differences should take into account local contextual information, we first apply a depthwise separable convolution to enhance local feature aggregation. This is followed by a Squeeze-and-Excitation block and a $1\times 1$ convolution layer, which are used to perform dynamic channel-wise weighting and adjust the number of output channels accordingly.
@@
After the above operations, we obtain the transformed feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, which are then fed into the Get Difference Attention module to generate the difference feature map $D_{i}$. This module is inspired by the core concept of attention mechanisms: measuring the similarity between vectors via dot products to assign corresponding token weights. Specifically, we first apply identical linear projection to $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, mapping them into a shared high-dimensional space and obtain the query Q and the key K. Next, we compute the dot product between corresponding vectors in Q and K to derive the initial attention mask M. Unlike standard attention mechanisms, our goal is to emphasize the features that exhibit greater differences. To this end, we invert the attention mask M, such that vector pairs with larger dot products (i.e., similar directions) receive lower weights, while those with smaller dot products (i.e., more distinct features) are assigned higher weights.This inversion encourages the model to focus more on significantly different features. Finally, we apply a Sigmoid function to normalize the inverted scores, resulting in the final attention mask $M^{\prime}$. To mitigate the risk of vanishing gradients caused by excessively large dot product values, a scaling factor $\sqrt{d_{k}}$(where $d_{k}$ represents the vector dimension) is applied during the dot product computation. This process can be expressed as:
$\textbf{Q}=Conv_{1\times 1}(f^{\prime}_{1\_i}),$
$\textbf{K}=Conv_{1\times 1}(f^{\prime}_{2\_i}),$
$\textbf{M}\supset m_{ij}=\frac{q_{ij}\cdot k_{ij}}{\sqrt{d_{k}}},\{i=1,2,\dots,W;j=1,2,\dots,H\},$
$^{\prime}=Sigmoid(-\textbf{M}).$ M ) .
@@
Finally, we compute teh value matrix V by calculating the Manhattan distance between the transformed feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$. The final output $D_{i}$ is then obtained by Hadamard product of the attention mask $\textbf{M}^{\prime}$ and value matrix V.
@@
$D_{i}=\textbf{M}^{\prime}\otimes\textbf{V}.$

### v2

@@ Abstract
Remote sensing change detection is essential for monitoring urban expansion, disaster assessment, and resource management, offering timely, accurate, and large-scale insights into dynamic landscape transformations. While deep learning has revolutionized change detection, the increasing complexity and computational demands of modern models have not necessarily translated into significant accuracy gains. Instead of following this trend, this study explores a more efficient approach, focusing on lightweight models that maintain high accuracy while minimizing resource consumption, which is an essential requirement for on-satellite processing. To this end, we propose FlickCD, which means quick flick then get great results, pushing the boundaries of the performance-resource trade-off. FlickCD introduces an Enhanced Difference Module (EDM) to amplify critical feature differences between temporal phases while suppressing irrelevant variations such as lighting and weather changes, thereby reducing computational costs in the subsequent change decoder. Additionally, the FlickCD decoder incorporates Local-Global Fusion Blocks, leveraging Shifted Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA) to effectively capture semantic information at multiple scales, preserving both coarse- and fine-grained changes. Extensive experiments on four benchmark datasets demonstrate that FlickCD reduces computational and storage overheads by more than an order of magnitude while achieving state-of-the-art (SOTA) performance or incurring only a minor (<1% F1) accuracy trade-off. The implementation code is publicly available at https://github.com/xulsh8/FlickCD.
Keywords Remote Sensing; Change Detection; Lightweight Deep Learning
@@ 1 Introduction
With the rapid advancement of remote sensing technologies, it has become possible to acquire an increasing number of high-resolution remote sensing images. These abundant and high-quality data sources have greatly enhanced the development of various remote sensing applications, among which change detection (CD) has particularly benefited as a key task. The target of CD is to identify the changes in the same region by analyzing multi-temporal remote sensing images acquired at different times. This technique can be applied to various fields, such as disaster damage assessment [1, 2], urban planning [3, 4], agricultural management [5] and environment monitoring [6]. The CD techniques make it possible to gain a more comprehensive understanding of the regional differences, enabling the prediction of future changes and the formulation of corresponding measures.
@@
With the growing success of deep learning in various computer vision tasks, it has also emerged as a powerful tool for CD. However, deep learning-based CD methods face two major challenges. The first challenge is how to better balance model performance and resource requirements. To achieve better feature extraction capabilities, change detection models often increase their depth or width a lot. For example, RCTNet [7] increases depth by stacking convolutional layers to fuse multi-scale feature maps, while ChangeMamba [8] expands the receptive field using state-space models to capture global semantic information. However, these operations require a substantial number of parameters and extensive computational resources, making it challenging to deploy them on resource-constrained devices, such as drones and satellites, which are commonly used for acquiring remote sensing images. Currently, several studies have also been conducted on lightweight change detection models [9, 10]. Nevertheless, achieving a lightweight design often results in performance degradation and leads to a reduced ability to extract difference information.
@@
The second challenge is how to ignore irrelevant differences while focusing on capturing significant changes. Irrelevant differences can arise due to various factors, including lighting variations, atmospheric conditions, and seasonal changes. To extract the critical changed information between two temporal images, various change detection models employ different approaches. Most models obtain difference information by subtracting the feature maps and taking the absolute value. [11, 9, 7]. Some models merge the two temporal feature maps along the channel dimension and use convolution operations to extract difference information [12, 10], while others perform layer-wise or multi-level feature fusion to achieve this [8, 13]. However, each of these methods has its own limitations: some of them fail to filter out irrelevant changes, making it difficult for the model to distinguish between insignificant differences and critical changes. Others require substantial computational resources and a large number of parameters to obtain difference information, thereby increasing the overall resource demand of the model.
@@
To address the aforementioned challenges, we propose FlickCD, a lightweight model that achieves an excellent balance between resource efficiency and performance. As shown in Fig. 1, we select several lightweight models and state-of-the-art models for comparison. It can be observed that while maintaining a lightweight design, FlickCD achieves an average performance close to that of state-of-the-art models.
FlickCD adopts an encoder-decoder architecture, with RepViT [14] chosen as its encoder. This encoder has the advantages of low latency and a small number of parameters. Unlike most change detection models, we utilize only the feature maps extracted from the first three stages of the encoder, as we observed that incorporating higher-level features does not lead to significant performance improvement, while substantially increasing the number of parameters and computational cost. After extracting feature maps from two temporal images using a shared-weight encoder, we employ the Enhanced Difference Module (EDM) to capture the difference information between them. The lightweight EDM can enhance the difference features while suppressing the irrelevant noise, thereby helping the decoder more accurately extract the changed features of interest. Then, we design the decoder using a bottom-up, layer-by-layer fusion approach. Before fusion, we incorporate a Local-Global Fusion Block (LGFB) to further integrate features. LGFB primarily utilizes attention mechanisms and consists of two components: Sliding-Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA). SWSA uses window-based attention to constrain focus to local regions, which helps the model better capture spatial relationships between nearby pixels and improves the accuracy of local object recognition. Subsequently, EGSA is employed to consider the relationship between pixels across the global scope, which is particularly critical when the changed regions consist of a lot of similar small areas. After multiscale extraction and fusion, the process yields a high-quality binary change map as the final result.
Through extensive experiments, we validate the effectiveness of the modules in FlickCD and demonstrate that FlickCD achieves a better trade-off and higher efficiency compared to other models. Our contributions can be summarized as follows:
@@
  • We conduct experiments to validate the efficiency of FlickCD among current CD models. It achieves low parameter count and computation cost while maintaining high performance, and even achieves state-of-the-art on some datasets.
@@ 2.1 Deep Learning-based Change Detection
With the remarkable success and high performance of deep learning techniques, the change detection (CD) community has increasingly adopted them to address the CD task. One of the earliest CNN-based approaches is the FC series [15], which introduced two fundamental strategies for handling dual-temporal inputs: (1) early fusion, where the two input images are first concatenated and then fed into the network; and (2) Siamese architectures, which process the two images with shared-weight encoders followed by a difference computation. These two strategies have laid the foundation for most subsequent CD models. After that, the emergence of Transformer architectures, which leverage the attention mechanisms, has brought new possibilities to CD. For instance, BIT [11] adopts a CNN backbone for encoding and decoding, while employing a Transformer to capture differences between the input images. ChangeFormer [12] further advances this direction by directly applying a shared-weight Transformer to handle the CD task. In addition to the conventional global self-attention, various attention mechanisms tailored for CD have also been explored. These include spatial and channel attention [16, 17, 18], as well as the cross-attention mechanism [19], which enhance the effectiveness and robustness of CD models.
@@ 2.2 Lightweight Strategies in Deep Learning
Striking a perfect balance between model efficiency and performance is always a critical research direction. To enable the development of deep learning models on resource-constrained devices, a variety of lightweight strategies have been proposed across different model architectures. In CNN-based strategies, Depth-wise Separable Convolution [20] extracts features by decoupling the spatial and channel dimensions. Group Convolution [21] divides channels into groups and performs convolution operations within each group independently. More recently, FasterNet [22] introduces Partial Convolution, which computes only a subset of channels in the feature maps. In Transformer-based strategies, they are primarily implemented in two ways. First, limit the attention computation scope to local windows or sparse regions [14, 23, 24]. Although this strategy may result in information loss, allocating limited parameters to more likely relevant tokens can lead to minimal performance degradation while reducing redundant information. Second, utilize weight-sharing mapping to obtain the attention matrix [23, 25]. Applying the same linear mapping to correlated regions reduces parameter count and projects them into the same high-dimensional space. There has also been research on lightweight models for CD tasks. USSFC-Net [9] uses dilated convolutions cyclically as the basic module, and constructs the Encoder and Decoder by stacking this module. TinyCD [10] employs the three-stage EfficientNet [26] as its backbone and utilizes group convolution to compute difference information. Although these networks have a low number of parameters, they suffer from high latency and lower performance.
@@ 3.1 Overview
The overview of the model structure is shown in the Fig. 2(a). At first, input the bi-temporal images, $\mathbf{T}_{1}$ and $\mathbf{T}_{2}$, both with dimensions $W\times H\times 3$, into the weight-sharing encoder. We selected the lightest version of RepViT [14] with pre-trained weights as the encoder. Then, we only use the feature maps from the first three stages of the encoder. On one hand, our experiments demonstrate that the feature map from the fourth stage does not improve the model’s performance. On the other hand, removing the fourth stage can reduce the number of parameters and computational complexity. This process can be expressed as:
$f_{1\_1},f_{1\_2},f_{1\_3}=Encoder(\mathbf{T}_{1}),$
$f_{2\_1},f_{2\_2},f_{2\_3}=Encoder(\mathbf{T}_{2}).$
The encoder extracts features hierarchically, where downsampling and channel doubling are applied at each stage. For the feature maps $f_{1\_i}$ and $f_{2\_i},i\in\{1,2,3\}$, the spatial resolution at each stage is $\frac{W}{2^{i+1}}\times\frac{H}{2^{i+1}}$. Then, we use the Enhanced Difference Module (EDM) to extract the difference feature between $f_{1}$ and $f_{2}$ at each stage. Additionally, to facilitate the multi-stage feature fusion in the decoder, the channel dimension of the feature maps from all three stages are unified. We formulate this process as:
$D_{i}=EDM_{i}(f_{1\_i},f_{2\_i}),i\in\{1,2,3\}.$
After that, we propose a bottom-up lightweight decoder, which mainly consists of the Local-Global Fusion Block (LGFB) and additional Efficient Global Self-attention (EGSA). Through layer-by-layer fusion and upsampling, it generates the final binary result map.
@@ 3.2 Enhanced Difference Module
How to extract the difference information between two temporal images effectively is one of the main challenges of the CD task. A common approach for extracting the difference is to directly subtract the two feature maps and take the absolute value. However, since this approach computes pixel-wise differences with equal weights across all pixels, differences caused by irrelevant factors are preserved alongside meaningful change information and passed to the decoder. This makes it difficult for the decoder to accurately identify critical differences, leading to a decrease in accuracy. To address this issue, we propose the Enhanced Difference Module (EDM), which dynamically assigns weights to the pixel-wise differences. This enables the model to filter out irrelevant differences and enhance the identification of critical change information.
@@
The detailed architecture of the EDM is illustrated in Fig. 2(b). For a given stage $EDM_{i}$, the inputs are the corresponding feature maps $f_{1\_i}$ and $f_{2\_i}$. Since the magnitude of pixel-wise differences should take local contextual information into account, we first apply a depthwise separable convolution to enhance local feature aggregation. This is followed by a Squeeze-and-Excitation block and a $1\times 1$ convolution layer, which are used to perform dynamic channel-wise weighting and adjust the number of output channels accordingly.
@@
After the above operations, we obtain the feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, which are then fed into the Difference-Aware Attention module to generate the difference feature map $D_{i}$. This module is inspired by the core concept of attention mechanisms: measuring the similarity between vectors via dot products to assign corresponding token weights. Specifically, we first apply identical linear projection to $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, mapping them into a shared high-dimensional space and obtain the query Q and the key K. Next, we compute the dot product between corresponding vectors in Q and K to derive the initial attention mask M. Unlike standard attention mechanisms, our goal is to emphasize the features that exhibit greater differences. To this end, we invert the attention mask M, such that vector pairs with larger dot products (i.e., similar directions) receive lower weights, while those with smaller dot products (i.e., more distinct features) are assigned higher weights. This inversion encourages the model to focus more on significantly different features. Finally, we apply a Sigmoid function to normalize the inverted scores, resulting in the final attention mask $\textbf{M}^{\prime}$. To mitigate the risk of vanishing gradients caused by excessively large dot product values, a scaling factor $\sqrt{d_{k}}$ (where $d_{k}$ represents the vector dimension) is applied during the dot product computation. This process can be expressed as:
$\textbf{Q}=Conv_{1\times 1}(f^{\prime}_{1\_i}),\textbf{K}=Conv_{1\times 1}(f^{\prime}_{2\_i}),$
$\textbf{M}\supset m_{ij}=\frac{q_{ij}\cdot k_{ij}}{\sqrt{d_{k}}},i\in\{1,2,\dots,W\};j\in\{1,2,\dots,H\},$
$\textbf{M}^{\prime}=Sigmoid(-\textbf{M}).$
@@
Finally, we compute the value matrix V by taking the element-wise absolute difference between the feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$. The final output $D_{i}$ is then obtained by the Hadamard product of the attention mask $\textbf{M}^{\prime}$ and value matrix V.
@@
$D_{i}=\textbf{M}^{\prime}\odot\textbf{V}.$